OpenStudy (anonymous):

if the net price of a stove is $900 and the trade discount rate is 40 percent the list price is

OpenStudy (anonymous):

60/100 = 900 / x 90000/60 = 1500

OpenStudy (anonymous):

thanks that really helped me